OpenStudy (ohohaye):

I'm told to Evaluate sin-1 1/2, but I don't understand what I'm supposed to do, please help me

OpenStudy (displayerror):

You're asked to evaluate \(\sin^{-1} \frac{1}{2}\). Unlike \(\sin \theta\), which gives you a measurement for a given angle \(\theta\), the inverse sin, \(\sin^{-1} a\), gives you an angle for a given measurement \(a\). In this case, the given measurement is 1/2. Now think: the \(\sin\) of what angle gives a measurement of 1/2? (Hint: Think of your standard triangles 45-45-90 and 30-60-90)

OpenStudy (ohohaye):

so am I supposed to put a number or something else?

OpenStudy (displayerror):

Do you remember the SOH CAH TOA acronym? \(\sin \theta\) is a ratio of the opposite side and hypotenuse of a right triangle for a given angle \(\theta\). Because you're given \(\sin^{-1} \frac{1}{2}\), you know that the opposite side has measurement 1 and the hypotenuse has measurement 2. The question is asking: the \(\sin\) of what angle gives you such a ratio?

OpenStudy (ohohaye):

oh ok, so I am supposed to figure out if is has a ratio of 1/2 or something like that?

OpenStudy (displayerror):

You're supposed to figure out the correct angle \(\theta\) that corresponds to a ratio of 1/2.

OpenStudy (ohohaye):

I don't undestand